RHIC pp Set-Up: Access for STAR and PHENIX went well, although the PHENIX Iris scanner failed again. The bo6-qd3-ps power supply was swapped during the access.
Yellow feedback ramp worked with efficiency of 99.5%. Lowered cavity voltage from 100 to 75kV and Landau phase was adjusted.
LINAC module 6 was arcing last night and had to be fixed. There was work on collimation during the LINAC repair.
The Yellow polarization was zero percent after feedback ramp. Work on Yellow polarization adjusting tunes and chromaticities. Measured 30% at store.. Filled with 56x56 store with 1e11 showing poor emittance and lifetime. Both improve after tune change. There was a Blue quench at 4:30AM caused by running the Blue Polarimeter target frame into the beam. This was accomplished by people using the pet pages to move the targets instead of the Polarimeter GUI. The users should NEVER use the pet pages to move the targets. Previously when the beam hit the target frame it destroyed all the targets on that frame. .
The main activity for today is to improve the Yellow polarization. Te question is how can the Blue polarization be 45% in Blue and only 30% in Yellow at injection? Polarization extracted from AGS is identical by definition. All the transfer elements in the AtR that effect the polarization are common to Blue and Yellow. So how can the polarization at injection be so different? Could it be a problem with the calibration of the Polarimeters? If Polarimeter is ready, measure polarization on the ramp.
Check Yellow snakes if Yellow and Blue polarization are different.
Scheduling Physicist: Mike Sivertz | Last updated: |
Email: sivertz@bnl.gov | |
Office Phone: 631-344-6102 | |
Digital Pager: 7528 or 4300 | |
Home Phone: 631-878-9020 |